# NoteLLM-2: Multimodal Large Representation Models for Recommendation

### v1

@@
Abstract
Large Language Models (LLMs) have demonstrated exceptional text understanding. Existing works explore their application in text embedding tasks. However, there are few works utilizing LLMs to assist multimodal representation tasks. In this work, we investigate the potential of LLMs to enhance multimodal representation in multimodal item-to-item (I2I) recommendations. One feasible method is the transfer of Multimodal Large Language Models (MLLMs) for representation tasks. However, pre-training MLLMs usually requires collecting high-quality, web-scale multimodal data, resulting in complex training procedures and high costs. This leads the community to rely heavily on open-source MLLMs, hindering customized training for representation scenarios. Therefore, we aim to design an end-to-end training method that customizes the integration of any existing LLMs and vision encoders to construct efficient multimodal representation models. Preliminary experiments show that fine-tuned LLMs in this end-to-end method tend to overlook image content. To overcome this challenge, we propose a novel training framework, NoteLLM-2, specifically designed for multimodal representation. We propose two ways to enhance the focus on visual information. The first method is based on the prompt viewpoint, which separates multimodal content into visual content and textual content. NoteLLM-2 adopts the multimodal In-Content Learning method to teach LLMs to focus on both modalities and aggregate key information. The second method is from the model architecture, utilizing a late fusion mechanism to directly fuse visual information into textual information. Extensive experiments have been conducted to validate the effectiveness of our method.
1 Introduction
Figure 1: Two strategies for using LLMs to enhance multimodal representations.
@@ 1 Introduction
With the development of the internet, most online platforms are full of multimodal information to drive user engagement [51, 50, 53]. Multimodal recommendation is a necessary service for these platforms [30], which recommends items based on users’ interest on item multimodal information. These techniques mainly rely on multimodal representations, extracting raw multimodal content into dense embeddings [61], and then measuring similarity through these multimodal embeddings. The enhancement of multimodal representations can fundamentally improve numerous downstream tasks, such as retrieval [31], clustering [34], and item modeling [60, 51].
@@
Many existing works have investigated multimodal representations, such as CLIP [37], METER [14] and Siglip [58]. Although great success has been achieved in these domains, the number of parameters in most multimodal representation models still possesses the large potential for scaling [45, 46]. However, existing works [45, 46] primarily focus on scaling vision models to improve multimodal representation, overlooking the textual branch. At the same time, due to the lack of sufficient pure text-oriented pre-training [37, 45, 46, 58], the text branch in multimodal representation models often exhibits suboptimal understanding of text information [6, 59]. Therefore, it is necessary to enhance text understanding in multimodal representations models.
@@
Recently, Large Language Models (LLMs) have shown impressive capabilities in text understanding and generating [18, 7, 2, 47]. Several works show the superior performance of LLMs in generating textual embeddings [34, 60, 19, 31, 4]. However, the use of LLMs for enhancing text understanding in multimodal representations is underexplored. As shown in Figure 1, one feasible method is to adapt Multimodal Large Language Models (MLLMs) [29, 25, 2, 22] for representation tasks. Nevertheless, pre-training MLLMs requires collecting high-quality, web-scale multimodal data, resulting in complex training procedures and high costs [25, 2]. This leads the community to rely heavily on open-source MLLMs, hindering customized training using the most advanced LLMs and vision encoders for representation scenarios. Therefore, we aim to design an end-to-end training method that customizes the integration of any existing LLMs and vision encoders to construct efficient multimodal representation models. This end-to-end method is practical when the target LLMs lack pre-trained visual components [11, 18, 1], as it only needs target domain data, eliminating pre-training data collection. Besides, we can optimize the customized model structure for representation tasks to achieve better efficiency while ensuring performance. We call these models for enhancing multimodal representation with LLMs as Multimodal Large Representation Models (MLRMs).
In this work, we explore MLRMs in item-to-item (I2I) multimodal recommendation scenarios. We conduct fine-tuned preliminary experiments to examine our goal. Our findings reveal that representations of naive fine-tuned end-to-end MLRMs are biased towards textual information, ignoring visual content. Visual information is mainly used to amplify the key element in multimodal content, rather than being directly aggregated by representations. To solve this challenge, we propose a novel training framework, called NoteLLM-2, for multimodal representation tasks. We design two methods to mitigate visual modality neglect. The first method is multimodal In-Context Learning (mICL). This method separates multimodal content into visual and textual components, subsequently compressing the content into two modality-compressed words. Each modality-compressed word conducts in-batch contrastive learning with its corresponding modality-compressed word. During inference, mICL can aggregate multimodal information in an In-Context Learning (ICL) way [12, 48]. The second method is the late fusion mechanism. Most MLLMs adopt early fusion, which maps visual features into the textual feature space of LLMs and interact visual features with text embeddings through LLMs [29, 25, 2, 22]. Even though early fusion can deepen the multimodal information interaction, it can easily lead to the loss of pure visual information during the multimodal fusion conducted by LLMs [25, 62]. We incorporate late fusion based on early fusion, retaining more visual information by delaying the fusion process. Specifically, we adopt a multimodal gate fusion mechanism [8], which takes visual representations from the visual encoder and representations from the LLM to generate final multimodal representations. We conduct extensive experiments to demonstrate our NoteLLM-2 effectiveness. Our paper makes the following contributions:
  • To the best of our knowledge, we are the first to explore the multimodal representation assisted by LLMs. We propose an end-to-end training method that customizes the integration of any existing LLMs and vision encoders to form efficient MLRMs.
  • We reveal that fine-tuning unaligned MLRMs for representation tasks can cause the modality imbalance, overlooking visual information. To overcome this problem, we design a novel framework, NoteLLM-2, which contains two mechanisms, mICL and late fusion.
  • Extensive experiments have demonstrated the effectiveness of our proposed framework for multimodal recommendation representation.
Our paper is organized as follows: We begin with a preliminary investigation of MLRMs in multimodal I2I recommendation scenarios and analyze the results. Then, inspired by this analysis, we design our NoteLLM-2. And we conduct experiments to demonstrate the effectiveness of our NoteLLM-2. Besides, related works and additional details are presented in Appendix A.
2 Preliminary Investigation
2.1 Preliminaries
Problem Statement. In our scenarios, every item represents a note, which is generated by users and expresses their life experiences. We show some notes in Appendix H. We assume $\mathcal{N}=\{n_{1},n_{2},...,n_{m}\}$ as note pool, where $m$ is the number of notes. Each note contains textual information, like the title, topic, and content, as well as the image. We denote the $i$-th note as $n_{i}=(t_{i},tp_{i},ct_{i},v_{i})$, where $t_{i}$, $tp_{i}$, $ct_{i}$, $v_{i}$ mean the title, the topic, the content and the image respectively. Given a query note $n_{i}$, the I2I recommendation system ranks the note pool $\mathcal{N}\backslash\{n_{i}\}$ according to the multimodal content similarity to the given note. The objective of this system is to prioritize the ranking of the related target note.
Dataset Construction. Following [60], we utilize the co-occurrence mechanism to build related note pairs based on user behaviors without human annotations. This mechanism assumes that notes frequently read together are likely related. Then, we count the co-occurrence score in which users viewed note $n_{A}$ and subsequently clicked on note $n_{B}$ as follows:
$s_{n_{A}\rightarrow n_{B}}=\sum_{u\in U_{n_{A}\rightarrow n_{B}}}\frac{1}{N_{u}},$ (1)
where $s_{n_{A}\rightarrow n_{B}}$ means the co-occurrence score from note $n_{A}$ to note $n_{B}$, $U_{n_{A}\rightarrow n_{B}}$ is the set of users who viewed note $n_{A}$ and subsequently clicked on note $n_{B}$, and $N_{u}$ denotes the quantity of the note set clicked by the user $u$ in the user behavior data. After calculating the co-occurrence score for all note pairs, we construct the set of co-occurrence scores $\mathcal{S}_{n_{i}}$ from note $n_{i}$ to all other notes. Specifically, $\mathcal{S}_{n_{i}}$ is defined as $\{{s_{n_{i}\rightarrow n_{j}}}|1\leq j\leq m,i\neq j\}$. Next, we remove outlier from $\mathcal{S}_{n_{i}}$ with co-occurrence scores above $up$ or below $low$. Finally, we select the $t$ notes with the highest co-occurrence scores from the filtered set as the related notes for note $n_{i}$.
@@ 2.1 Preliminaries
Figure 2: The frameworks of MLRMs. (a) is the basic method for representations. (b) is our representation method, which contains two easy and effective mechanisms to enhance multimodal representation ability, mICL and late fusion.
@@
Basic Representation Method of MLRMs. Following [19], we adopt a prompt with explicit one word limitation to compress the multimodal content into one embedding. Specifically, we use json format to show note content and our prompt is the following: {mdframed} Note Compression Prompt
@@
Note content: {’image’: <IMG>, ’title’: $t_{i}$ , ’topic’: $tp_{i}$ , ’content’: $ct_{i}$}. Compress this note into one word:" In this template, <IMG> is a placeholder, which will be replaced by vision embeddings processed from raw images $v_{i}$.
@@
The overall framework is shown in Figure 2 (a). This framework first utilizes the vision encoder $V_{\theta}$ to extract the pre-processed image $v_{i}$ into visual features $Z_{v}\in\mathbb{R}^{L\times h_{v}}$ in the notes $n_{i}$, where $L$ is the length of visual features, $h_{v}$ is the dimension of visual features and $Z_{v}=V_{\theta}(v_{i})$. Then, the connector $C_{\pi}$ transforms the visual features $Z_{v}$ into the word embedding space of LLMs to form visual embeddings $E_{v}\in\mathbb{R}^{L_{c}\times h_{t}}$, where $L_{c}$ is the length of visual embeddings input into LLMs, $h_{t}$ is the dimension of LLMs’ hidden states and $E_{v}=C_{\pi}(Z_{v})$. After processing images, the text prompts are tokenized into discrete indexes, then form the text word embedding $E_{t}\in\mathbb{R}^{T\times h_{t}}$, where $T$ is the length of text tokens. To insert the visual embeddings into the correct position in text embeddings, we replace word embeddings in positions of the <IMG> token by the visual embeddings $E_{v}$, forming the multimodal embeddings $E_{m}\in\mathbb{R}^{(L_{c}+T-1)\times h_{t}}$. Last, we use LLMs $LLM_{\mu}$ to process the multimodal embeddings to generate last hidden states $H\in\mathbb{R}^{(L_{c}+T-1)\times h_{t}}$, where $H=LLM_{\mu}(E_{m})$. We take last embeddings of $H$ as the note representation $\boldsymbol{n}_{i}$ for note $n_{i}$. This method constrains LLMs to compress the multimodal content into one embedding in the form of the next token prediction.
@@
However, LLMs are trained using the language modeling loss [29, 25, 3, 22], which significantly differs from representation tasks. To bridge this gap, we employ contrastive learning, a prevalent technique for training embeddings [34, 60, 19, 31]. Specifically, each minibatch contains $B$ related note pairs $(n_{i},n_{i}^{+})$, resulting in a total of $2B$ notes per minibatch. Following [36], we use gradient descent to minimize the contrastive loss as follows:
$\mathcal{L}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{i}^{+})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{j})\cdot e^{\tau}}},$ (2)
@@
where $\tau$ means the learnable temperature and $sim(a,b)=a^{\top}b/(\|a\|\|b\|)$. $\mathcal{L}(\pi,\mu)$ means that we only update the connector $C_{\pi}$ and LLMs $LLM_{\mu}$, while keep the vision encoder $V_{\theta}$ frozen, aiming for the larger batch size and better performance [22].
2.2 Datasets and Experimental Settings
Datasets. We obtain a real-world multimodal I2I dataset from our platform. We create the training dataset by randomly choosing related note pairs from user behaviour data gathered over two weeks. One-tenth of pairs in this training dataset is used for the validation set. Next, we randomly select notes from the upcoming week to form the note pool of testing set, excluding any notes that are already in the training dataset. Then, we count the related note pairs in the note pool of testing set. The training dataset contains $1.5M$ notes and $1.1M$ note pairs. The testing dataset contains $0.5M$ notes and $21K$ note pairs. We provide detailed statistics of the training and testing dataset in Appendix C.
To better evaluate the multimodal representation ability of MLRMs rather than relying heavily on the text modality, we collect pairs that include short notes in the test dataset. We define notes with a token length of less than $50$ as short notes, which approximately constitute $10\%$ of the total test notes. We categorize pairs with short query notes as short query pairs and those with short target notes as short target pairs. In the test dataset, the number of short query pairs is $5,620$ and the number of short target pairs is $5,582$.
Experimental settings. In related note pair construction, we set the upper bound of the co-occurrence score $up$ as $30$ and the lower bound $low$ as $0.01$. And we set $t$ as $3$. To maintain context length restrictions, we truncate the titles exceeding $20$ tokens, and truncate the contents exceeding $80$ tokens. In the fine-tuned experiments, for fair comparison, we add a linear projector to diminish the dimension of note embeddings to $64$. The batch size $B$ is set to 128128128128111We train models on 8 $\times$ 80GB Nvidia A100 GPUs and the batch size per each GPU is $16$.. Each batch contains $256$ notes. The temperature $\tau$ is initialized as $3$. We provide more training details in Appendix D.
To evaluate the representation ability of MLRMs, we rank all notes in the note pool, excluding the query note, according to the content of query note. We compute recall scores based on the position of the target note in the ranked list. We report Recall@100, Recall@1k and Recall@10k on all pairs, short query pairs and short target pairs in the test dataset.
Table 1: Representation performance of different models (%).
@@ 2.2 Datasets and Experimental Settings
MQwen-bigG Multimodal 77.64 92.89 98.91 57.45 80.92 95.73 57.90 83.49 96.99
@@
Qwen-VL-Chat [^2] Due to Qwen-VL-Chat having $256$ visual embeddings per image, this model is trained on 32 $\times$ 80GB Nvidia A100 GPUs and the batch size per each GPU is $4$. Multimodal 78.54 93.77 99.03 60.43 83.26 96.60 61.63 85.48 97.89
2.3 Performance of Multimodal Representation of Fine-tuned MLRMs

### v2

@@
Abstract.
Large Language Models (LLMs) have demonstrated exceptional proficiency in text understanding and embedding tasks. However, their potential in multimodal representation, particularly for item-to-item (I2I) recommendations, remains underexplored. While leveraging existing Multimodal Large Language Models (MLLMs) for such tasks is promising, challenges arise due to their delayed release compared to corresponding LLMs and the inefficiency in representation tasks. To address these issues, we propose an end-to-end fine-tuning method that customizes the integration of any existing LLMs and vision encoders for efficient multimodal representation. Preliminary experiments revealed that fine-tuned LLMs often neglect image content. To counteract this, we propose NoteLLM-2, a novel framework that enhances visual information. Specifically, we propose two approaches: first, a prompt-based method that segregates visual and textual content, employing a multimodal In-Context Learning strategy to balance focus across modalities; second, a late fusion technique that directly integrates visual information into the final representations. Extensive experiments, both online and offline, demonstrate the effectiveness of our approach. Code is available at https://github.com/Applied-Machine-Learning-Lab/NoteLLM.
Multimodal Large Language Model; Recommendation; Multimodal Representation
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709440isbn: 979-8-4007-1245-6/25/08ccs: Information systems Recommender systems
1. Introduction
Figure 1. Two strategies for using LLMs to enhance multimodal representations. The first strategy transfers existing pre-trained MLLMs into representation tasks. The second strategy fine-tunes the integration of LLMs and vision encoders end-to-end, requiring no alignment and offering better practicality and efficiency.
@@ 1. Introduction
With the development of the internet, most online platforms are full of multimodal information to drive user engagement (Wu et al., 2022; Wei et al., 2019; Xun et al., 2021). Multimodal recommendation is a necessary service for these platforms (Liu et al., 2023a), which recommends items based on users’ interest in item multimodal information. These techniques mainly rely on multimodal representations, extracting raw multimodal content into dense embeddings (Zhang et al., 2020), and then measuring similarity through these multimodal embeddings. The enhancement of multimodal representations can fundamentally improve numerous downstream tasks, such as retrieval (Ma et al., 2023), clustering (Muennighoff et al., 2024), and item modeling (Wu et al., 2022).
@@
Many existing efforts have investigated multimodal representations, such as CLIP (Radford et al., 2021) and METER (Dou et al., 2022). Despite significant success in these areas, the number of parameters in most multimodal representation models still has great potential for scaling (Sun et al., 2023, 2024). However, existing works (Sun et al., 2023, 2024) primarily focus on scaling vision models to improve multimodal representation, overlooking the textual branch. At the same time, due to the lack of sufficient pure text-oriented pre-training (Radford et al., 2021; Sun et al., 2023, 2024; Zhai et al., 2023), the text branch in multimodal representation models often exhibits suboptimal understanding of text information (Chen et al., 2023; Zhang et al., 2024c). Therefore, it is necessary to enhance text understanding in multimodal representation models.
@@
Recently, Large Language Models (LLMs) have shown impressive capabilities in text understanding and generating (Jiang et al., 2023b; Chiang et al., 2023; Bai et al., 2023a; Touvron et al., 2023). Several works show the superior performance of LLMs in generating textual embeddings (Muennighoff et al., 2024; Zhang et al., 2024a; Jiang et al., 2023a; Ma et al., 2023; BehnamGhader et al., 2024). However, the use of LLMs for enhancing text understanding in multimodal representations is underexplored. As shown in Figure 1, one feasible method is to adapt existing open-source Multimodal Large Language Models (MLLMs) (Liu et al., 2023b; Li et al., 2023; Bai et al., 2023a; Karamcheti et al., 2024) for representation tasks. However, developing MLLMs requires substantial high-quality multimodal data and significant computational resources for pre-training (Li et al., 2023; Bai et al., 2023a), resulting in the release of LLMs without multimodal capabilities ahead of their multimodal counterparts (DeepSeek-AI, 2024; Jiang et al., 2023b; lla, [n. d.]). This delay constrains customized training using the most advanced LLMs and vision encoders. Moreover, existing MLLMs are inefficient for representation scenarios as they primarily focus on multimodal understanding and generation (Shang et al., 2024). Therefore, we design an end-to-end fine-tuning method that customizes the integration of any existing LLMs and vision encoders to construct efficient multimodal representation models. This end-to-end method is practical when the target LLMs lack pre-trained visual components (DeepSeek-AI, 2024; Jiang et al., 2023b; lla, [n. d.]), as it only needs target domain data, eliminating pre-training data collection. Besides, we can optimize the customized model structure for representation tasks to achieve better efficiency while ensuring performance. We refer to these models for enhancing multimodal representation with LLMs as Multimodal Large Representation Models (MLRMs).
@@
Figure 2. The frameworks of MLRMs. (a) is the basic method for representations. (b) is our representation method, which contains two easy and effective mechanisms to enhance multimodal representation ability, mICL and late fusion.
In this work, we explore MLRMs in item-to-item (I2I) multimodal recommendation scenarios. We conduct preliminary fine-tuning experiments to examine our goal. Our findings reveal that naive end-to-end fine-tuned MLRMs are biased towards text, ignoring visual content. Images are mainly used to amplify key elements in multimodal content, rather than being directly aggregated by representations. To solve this challenge, we propose NoteLLM-2, a novel fine-tuning framework for multimodal tasks, with two methods to mitigate visual neglect. The first method is multimodal In-Context Learning (mICL). This method separates multimodal content into visual and textual components, subsequently compressing the content into two modality-compressed words. Each modality-compressed word conducts in-batch contrastive learning with its corresponding modality-compressed word. The second method is the late fusion mechanism. Most MLLMs adopt early fusion, which maps visual features into the textual feature space of LLMs and interacts visual features with text embeddings through LLMs (Liu et al., 2023b; Li et al., 2023; Bai et al., 2023a; Karamcheti et al., 2024). Even though early fusion can deepen the multimodal information interaction, it can easily lead to the loss of pure visual information during the multimodal fusion conducted by LLMs (Li et al., 2023; Zhang et al., 2024b). We incorporate late fusion based on early fusion, retaining more visual information by delaying the fusion process. Specifically, we adopt a multimodal gate fusion mechanism, which takes visual representations from the visual encoder and representations from the LLM to generate final multimodal representations. We conduct extensive experiments to demonstrate our NoteLLM-2 effectiveness. Our paper makes the following contributions:
  • To the best of our knowledge, we are the first to explore the multimodal representation assisted by LLMs in recommendation scenarios. We propose an end-to-end fine-tuning method that customizes the integration of any existing LLMs and vision encoders to form efficient MLRMs.
  • We reveal that fine-tuning unaligned MLRMs for representation tasks can cause modality imbalance by overlooking visual information. To overcome this problem, we design a novel framework, NoteLLM-2, which contains mICL and late fusion.
  • Extensive online and offline experiments have demonstrated the effectiveness of our proposed framework for multimodal recommendation representation.
2. Preliminary Investigation
2.1. Problem Statement
In our scenarios, every item represents a note, which is generated by users and expresses their life experiences. We assume $\mathcal{N}=\{n_{1},n_{2},...,n_{m}\}$ as note pool, where $m$ is the number of notes. Each note contains textual information, like the title, topic, and content, as well as the image. We denote the $i$-th note as $n_{i}=(t_{i},tp_{i},ct_{i},v_{i})$, where $t_{i}$, $tp_{i}$, $ct_{i}$, $v_{i}$ mean the title, the topic, the content and the image respectively. Given a query note $n_{i}$, the I2I recommendation system ranks the note pool $\mathcal{N}\backslash\{n_{i}\}$ according to the multimodal content similarity to the given note. The objective of this system is to prioritize the ranking of the related target note.
2.2. Basic Representation Method of MLRMs
@@
Following (Jiang et al., 2023a), we adopt a prompt with an explicit one-word limitation to compress the multimodal content into one embedding. Specifically, we use json format to show note content and our prompt is the following: {mdframed} Note Compression Prompt
@@
Note content: {‘image’: ¡IMG¿, ‘title’: $t_{i}$ , ‘topic’: $tp_{i}$ , ‘content’: $ct_{i}$}. Compress this note into one word:“ In this template, ¡IMG¿ is a placeholder, which will be replaced by vision embeddings processed from raw images $v_{i}$.
@@
The overall framework is shown in Figure 2 (a). This framework first utilizes the vision encoder $V_{\theta}$ to extract the pre-processed image $v_{i}$ into visual features $Z_{v}\in\mathbb{R}^{L\times h_{v}}$ in the notes $n_{i}$, where $L$ is the length of visual features, $h_{v}$ is the dimension of visual features and $Z_{v}=V_{\theta}(v_{i})$. Then, the connector $C_{\pi}$ transforms the visual features $Z_{v}$ into the word embedding space of LLMs to form visual embeddings $E_{v}\in\mathbb{R}^{L_{c}\times h_{t}}$, where $L_{c}$ is the length of visual embeddings input into LLMs, $h_{t}$ is the dimension of LLMs’ hidden states and $E_{v}=C_{\pi}(Z_{v})$. After processing images, the text prompts are tokenized into discrete indexes, then form the text word embedding $E_{t}\in\mathbb{R}^{T\times h_{t}}$, where $T$ is the length of text tokens. To insert the visual embeddings into the correct position in text embeddings, we replace word embeddings in positions of the ¡IMG¿ token by the visual embeddings $E_{v}$, forming the multimodal embeddings $E_{m}\in\mathbb{R}^{(L_{c}+T-1)\times h_{t}}$. Last, we use LLMs $LLM_{\mu}$ to process the multimodal embeddings to generate last hidden states $H\in\mathbb{R}^{(L_{c}+T-1)\times h_{t}}$, where $H=LLM_{\mu}(E_{m})$. We take the last embeddings of $H$ as the note representation $\boldsymbol{n}_{i}$ for note $n_{i}$. This method constrains LLMs to compress the multimodal content into one embedding in the form of the next token prediction.
@@
However, LLMs are trained using the language modeling loss (Liu et al., 2023b; Li et al., 2023; Bai et al., 2023b; Karamcheti et al., 2024), which significantly differs from representation tasks. To bridge this gap, we employ contrastive learning, a prevalent technique for training embeddings (Muennighoff et al., 2024; Zhang et al., 2024a; Jiang et al., 2023a; Ma et al., 2023). Specifically, each minibatch contains $B$ related note pairs, resulting in a total of $2B$ notes per minibatch. For any note $n_{i}$ (where $1\leq i\leq 2B$) in the minibatch, we denote its related note in the same minibatch as $n_{i}^{+}$. Following (Neelakantan et al., 2022), we use gradient descent to minimize the contrastive loss as follows:
(1) $\mathcal{L}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{i}^{+})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{j})\cdot e^{\tau}}},$
@@
where $\tau$ means the learnable temperature and $sim(a,b)=a^{\top}b/(\|a\|\|b\|)$. $\mathcal{L}(\pi,\mu)$ means that we only update the connector $C_{\pi}$ and LLMs $LLM_{\mu}$, while keep the vision encoder $V_{\theta}$ frozen, aiming for the larger batch size and better performance (Karamcheti et al., 2024).
Table 1. Detailed statistics of training and testing dataset.
                                  training dataset
  # notes 1,487,613 # note pairs 1,130,906
avg. # words per title 12.38 avg. # topic per note 5.26
avg. # words per topic 4.53 avg. # words per content 133.20
                                  testing dataset
  # notes 534,144 # note pairs 21,262
avg. # words per title 11.68 avg. # topic per note 4.58
avg. # words per topic 4.50 avg. # words per content 116.00
Table 2. Performance of different fine-tuned models (%). “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline.
@@
MQwen-bigG Multimodal 77.63 92.89 98.91 57.45 80.92 95.73 57.95 83.50 96.99
@@
Qwen-VL-Chat [^1] Due to Qwen-VL-Chat having $256$ visual embeddings per image, this model is trained on 32 $\times$ 80GB Nvidia A100 GPUs and the batch size per GPU is $4$. Multimodal 78.53* 93.76* 99.03* 60.41* 83.27* 96.60* 61.60* 85.48* 97.89*
2.3. Datasets and Experimental Settings
Dataset Construction Method. To construct pairs of related notes without relying on human annotations, we leverage the co-occurrence mechanism (Zhang et al., 2024a), which measures the relevance between notes by how often they are read together. We count the co-occurrence score in which users viewed note $n_{A}$ and subsequently clicked on note $n_{B}$ as follows:
(2) $s_{n_{A}\rightarrow n_{B}}=\sum_{u\in U_{n_{A}\rightarrow n_{B}}}\frac{1}{N_{u}},$
where $s_{n_{A}\rightarrow n_{B}}$ means the co-occurrence score from note $n_{A}$ to note $n_{B}$, $U_{n_{A}\rightarrow n_{B}}$ is the set of users who viewed note $n_{A}$ and subsequently clicked on note $n_{B}$, and $N_{u}$ denotes the number of distinct notes clicked by the user $u$ in the user behavior data. We compute the co-occurrence score for each note pair. Then, we construct the set $\mathcal{S}_{n_{i}}$, which contains co-occurrence scores from note $n_{i}$ to all other notes, i.e., $\mathcal{S}_{n_{i}}=\{{s_{n_{i}\rightarrow n_{j}}}|1\leq j\leq m,i\neq j\}$. Next, we remove outlier from $\mathcal{S}_{n_{i}}$ with co-occurrence scores above $up$ or below $low$. Finally, the top $t$ notes with the highest co-occurrence scores in the filtered set are viewed as the related notes for note $n_{i}$.
Datasets Details. We obtain a real-world multimodal I2I dataset from our platform. We create the training dataset by randomly choosing related note pairs from user behavior data gathered over two weeks. One-tenth of the pairs in this training dataset are used for the validation set. Next, the test note pool is collected by randomly selecting notes from the following week, while avoiding duplication with the training dataset. Then, we count the related note pairs in the test note pool. We provide detailed statistics of the training and testing dataset in Table 1.
To better evaluate the multimodal representation ability of MLRMs rather than relying heavily on the text, we collect pairs that include short notes in the test dataset. We define notes with a token length of less than $50$ as short notes, which approximately constitute $10\%$ of the total test notes. We categorize pairs with short query notes as short query pairs and those with short target notes as short target pairs. In the test dataset, the number of short query pairs is $5,620$ and the number of short target pairs is $5,582$.
Experimental settings. When constructing pairs of related notes, the upper bound of the co-occurrence score $up$ is $30$ and the lower bound $low$ is $0.01$. We set $t$ as $3$. We truncate titles longer than 20 words and content longer than 80 words to comply with inference length restrictions. In the fine-tuned experiments, for fair comparison, we add a linear projector to diminish the dimension of note embeddings to $64$. The batch size $B$ is set to 128128128128111Models trained on 8 $\times$ 80GB Nvidia A100 GPUs and the batch size per GPU is $16$.. There are $256$ notes in each batch. We start with a temperature $\tau$ set as $3$. We provide more training details in Appendix A.
To evaluate the representation ability for recall tasks, we create a similarity ranked list of all remaining notes in the test pool based on the query note’s content. We compute recall scores based on the position of the target note in the ranked list. We report Recall@100, Recall@1k, and Recall@10k on all pairs, short query pairs, and short target pairs in the test dataset. For robustness, we report the average results tested three times using random seeds {42,43,44}.
2.4. Performance of Multimodal Representation of Fine-tuned MLRMs